# LAVA: Explainability for Unsupervised Latent Embeddings

### v1

@@ 3 Experiments
We perform experiments to verify and examine the explanations of LAVA. We first look at a latent space as represented by LAVA localities, and show how different local subpatterns of correlation can exhibit different levels of similarity across the latent space. Then, we look at how module extraction reveals and utilizes these similarities by extracting modules of reoccurring subpatterns, and verify their properties mentioned in Section 2. Finally, we use LAVA modules to explain a UMAP embedding of a kidney single-cell gene expression dataset, revealing biologically relevant patterns.
@@
Datasets Our experiments are based on two datasets: MNIST handwritten digit images [39], and KPMP [^1] Kidney Precision Medicine Project: https://www.kpmp.org/. single-cell kidney gene expression [40]. The KPMP dataset measures the activity of $\sim$30k genes (features) for $\sim$225k kidney cells (samples) [41]. Experimental setup. We use the following UMAP embeddings [29]: for MNIST, we generate a UMAP embedding using default hyperparameters, while for the KPMP data we use the UMAP embedding provided with the dataset. For KPMP, we additionally remove mitochondrial and ribosomal genes and use only the top $1000$ most variable genes following standard single-cell analysis guidelines [42] (Appendix B.2). We apply LAVA to each dataset, with the neighborhood size hyperparameter $n$ set to 3000 for MNIST and $500$ for KPMP (Appendix A.1). For MNIST, we visualize the correlations of localities or modules as image-sized heatmaps, where the color of each pixel is determined by the sum of its correlations, and additional blue lines are drawn between correlated pixels (visible in most figures; details in Appendix B.4).
3.1 Local patterns, global subpatterns
With this experiment, we explore the similarity between correlation patterns across a latent space. Setup. We use LAVA locality representations of the MNIST UMAP, without extracted modules. As examples, we select two different localities, and then calculate cosine similarity between their and other localities’ respective representations. This is done using correlations for all feature pairs, or using subsets involving only features from the top or bottom half of the images. Intra- and intercluster heterogeneity. We observe that patterns of correlations between the original features can be highly dissimilar across localities, even within the same cluster of digits (first column of Figure 3). This can be conceptualized as a multivariate version of Simpson’s paradox, wherein different groupings of data – here LAVA localities – can show different trends in association [43], and highlights the fact that local feature associations cannot be trivially summarized across an entire dataset. Subpatterns of correlations alter similarity. Considering only locality correlations between pixels from the top or bottom half of the images, we observe variations in the cosine similarities across localities (second and third columns of Figure 3). This indicates how different localities across the latent space might share specific subpatterns of correlation, but not others, which is precisely what the module extraction in LAVA aims to capture.
Figure 3: Cosine similarities between each of two reference localities and all other localities of the latent MNIST UMAP space in three settings, with example visualizations of similar localities.
3.2 Module extraction quality and module properties
We analyze the quality of module extraction and the properties of LAVA modules for both datasets.
Figure 4: Module extraction results for MNIST and KPMP.
Setup. We run module extraction $R=10$ times with different numbers of modules $M$ for each dataset, reporting the main loss term (as cosine similarity instead of distance), overestimation ratio, and silhouette width score (Figure 4). Means and standard deviations of the loss and overestimation ratio are calculated based on the $10$ runs. Stability and reconstruction across runs. As expected, the loss term cosine similarity increases with larger numbers of modules, resulting in better reconstruction of the original correlations. For MNIST, we observe relatively low silhouette widths irrespective of the number of modules [44], and smaller standard deviations of the loss, pointing to successful optimization and thus a larger number of equally valid solutions. We hypothesize this is due to a large number of common correlation subpatterns, resulting in more ways to reconstruct locality correlations. For KPMP, we see a gradual decrease in stability (silhouette) with larger numbers of modules, indicating more stable solutions for smaller numbers, but similar instability with larger numbers. Number of modules. We decided to use $9$ modules for both datasets based on the criteria outlined in Subsection 2.3.3. Out of the $10$ runs, we chose the one with the lowest loss.
Modules reveal global subpatterns. For both datasets, the extracted modules represent globally reoccurring subpatterns not tied to neighboring localities (Figure 5). For MNIST, the first two modules (ordered by sum of presence in the latent space) capture common patterns of correlations between digits $3$, 55, and $8$, and between digits $4$, 77, and $9$ (Figure 5a). Visualizations of the two modules also show how the extracted correlations come from pixels shared across the digits these localities represent, as we would expect. We observe similar trends for KPMP embeddings (Figure 5b), with some modules present across most major cell types (module $3$), and others pertaining to more local regions of embeddings of more specific cell types (epithelial cells in the case of modules $1$, 22, and $4$). The module extraction in LAVA uses no information about the spatial organization of localities, allowing it to identify shared subpatterns across clusters, and differentiate patterns within clusters.
@@ 3.2 Module extraction quality and module properties
Figure 5: Labels and example module presences for the MNIST and KPMP UMAP embeddings.
Modules identify reoccurring, not averaged subpatterns. We observe that AMF extracts modules with overestimation error ratios of around $10\%$ or slightly higher (Figure 4). Errors made up of predominantly underestimation indicate that averaged patterns are avoided in favor of sparser, but actually present subpatterns, which can be seen as a type of quantile regression [37]. Combined with the semantics of cosine similarity, the presence of a module in two localities implies not only a similar subpattern of correlations, but also that the subpattern can directly be observed in both localities. We visualize this for MNIST by comparing two extracted modules to what they would look like if their LAVA extracted correlations were replaced by the presence-weighted average across localities.
Figure 6: Extracted modules vs. presence-weighted averages.
Multiple modules can coexist in a single locality. To investigate if LAVA modules can be co-present in the same locality, we scale the presences for each locality to sum to $1$, and calculate Shannon entropy to quantify module co-presence per locality in each dataset. Ignoring localities with summed presences <0.5<0.5 to prevent inflated entropies from scaling, we get average values of $0.68$ ($\pm 0.41$) and $1.41$ ($\pm 0.75$) for MNIST and KPMP, respectively (Figures 7a and 7b). For reference, 0 means no overlap between modules and $3.17$ corresponds to all $9$ modules being equally present. The obtained entropies indicate some overlap in both datasets, more so in KPMP. Figure 7c shows an example of how modules overlap to reconstruct correlations of a locality in MNIST. Allowing overlap, a given locality can be related to different subsets of localities through presences of modules, even when those modules represent partially overlapping subpatterns of correlation.
@@
Figure 7: Entropy of module presences per locality for the MNIST and KPMP datasets, with an MNIST example of how modules are combined to reconstruct localities.
3.3 Explaining embeddings of gene expression data (KPMP)
We investigate if LAVA modules can reveal biologically relevant patterns using the UMAP of KPMP data. Setup. The KPMP dataset includes metadata about the human donors of the cells: we look at disease state, which is either healthy, acute kidney failure (ACF), or chronic kidney disease (CKD). We first calculate percentages of samples from donor disease state per locality. We then calculate Pearson correlations and respective p-values between module presences and these percentages (ignoring localities where a module is not present) to gauge the association between module and disease state. We further investigate if genes in those modules are related to biological pathways related to diseased state (Appendix B.6). Module correlates with diseased samples. Cells from patients with different disease states are spread throughout the latent space, with no apparent clustering. However, the presence of module $4$ (Figure 5b) is positively correlated with the percentage of diseased samples ($0.23$, $p<0.01$), and of CKD samples specifically ($0.28$, $p<0.01$). Together with the fact that module $4$ is present across epithelial cells and not blood cells, this suggests that the correlation subpattern of module $4$ could be biologically relevant for kidney disease related pathways. Module genes and disease related pathways. Overrepresentation analysis of biological pathways [45] for the set of genes in module $4$ correlations exhibit $21$ significant pathways ($p<0.01$). To highlight biological relevance, we single out the more general “cellular response to stress” and the DNA damage response pathways, linked to CKD in previous studies [46, 47, 48, 49]. Overall, our findings suggest that the local organization of KPMP latent UMAP embeddings across kidney tissue cell types exhibits patterns of gene correlation related to kidney disease. This local structure of sample organization is not apparent in the clustering of the data, but notably discoverable through the LAVA modules.
@@ 3.3 Explaining embeddings of gene expression data (KPMP)
Figure 8: Module 4 for KPMP: overlap with chronic kidney disease and gene analysis.
4 Limitations
Defining neighborhood size. A larger limitation of LAVA is the definition of a relevant size for localities. Explanations based on local sample groupings are relevant for embeddings produced by techniques like non-linear manifold learning, which themselves try to preserve local relationships, but the choice of neighborhood size remains a challenge. The issue of defining meaningful localities relates more broadly to that of latent axes not corresponding to known quantities or concepts. While in unsupervised settings, exploration of multiple sizes might be of interest, in supervised settings or using metadata, size of locality could be motivated by changes in a target of interest. Association as correlation, and directionality. We estimate pairwise association between original features with Spearman correlations, therefore more complex associations will not be captured. Furthermore, our module extraction assumes absolute values, and cannot be trivially extended to differentiate between negative and positive correlation because of its use of the maximum function. This means LAVA cannot differentiate correlation reversals. Module extraction. As we saw in MNIST experiments, module extraction can lead to multiple equally good solutions. Assuming the problem setting is well-defined, each of these solutions is worth exploring, which might not always be practically feasible. We did not look into whether some modules are consistent across the different runs, which might be an exploitable finding. Choosing the number of modules used based on a stability/accuracy trade-off also retains some ambiguity, but is inherent to the unsupervised setting. Unstructured data. While we experiment with MNIST, our approach is generally suitable for tabular, as opposed to unstructured data such as images and text. With larger adaptations, a concept-based approach looking at presences of concepts rather than correlations between individual features, might be applicable [50].
@@ 5 Conclusions
In this work, we framed the problem of latent space explainability as enabling reasoning about the relative spatial organization of embeddings and their original data. We focused on explainability for latent embeddings generated by manifold learning that only produce embeddings and no mapping function. To that effect, we introduced the LAVA method, which describes local groups of data in the latent space through observed correlations between features in the original data, and relates them globally to one another by revealing reoccurring subpatterns of correlation throughout the latent space. In future work, we envision further exploring the idea of reasoning about latent embeddings spatially, especially for higher-dimensional latent spaces and supervised tasks, in which the spatial relations between embedded samples can be put into context, and differentiated from one another based on specific targets of interest, which we anticipate to be possible building on top of existing feature and sample importance methods. We see this approach having potential to extract more insight from ML models than existing methods, further aiding in model development and knowledge discovery.
Acknowledgments and Disclosure of Funding
@@
Funded by the National Institute of Diabetes and Digestive and Kidney Diseases (Grant numbers: U01DK133081, U01DK133091, U01DK133092, U01DK133093, U01DK133095, U01DK133097, U01DK114866, U01DK114908, U01DK133090, U01DK133113, U01DK133766, U01DK133768, U01DK114907, U01DK114920, U01DK114923, U01DK114933, U24DK114886, UH3DK114926, UH3DK114861, UH3DK114915, UH3DK114937)
@@ A.1 Locality Definition
A.1.1 Hyperparameter $n$: size of locality
@@
The choice of $n$ is important as the definition of localities has a substantial influence on correlations that LAVA will calculate in the step (2), and subsequently on module extraction in step (3). A data-driven way of choosing the value $n$ can be looking at which size of latent neighborhood most resembles the neighborhoods in the original feature size. In choosing such a $n$, we attempt to capture localities at a scale that most accurately represents organization in the original feature space. Another, more targeted, approach is motivating the choice of $n$ using metadata of interest. This might ground the scale of localities to existing knowledge and make findings more relevant for a downstream task. We employ both approaches in our experiments in Section 3, and deliberate further in B.3.
A.1.2 Defining the number of localities
$\centering\ell=[E*o/n],\@add@centering$ (8)
where [][\,] denotes rounding to the nearest integer. The hyperparameter $o$ determines the number of localities given a dataset and neighborhood size $n$. If all samples were of a similar centrality (e.g., if data are uniformly spread in some space), $o=x$ would mean that each sample would be represented by approximately $x$ localities. A larger $o$ will produce more overlap between localities and enable a representation that more closely approximates the samples’ original in-degree centralities, but will be more expensive computationally and harder to manage in subsequent analysis. We recommend using smaller values of $o$ in the range of $2$ to $20$, so that the resulting number of localities $\ell$ is manageable, but still captures the various correlation patterns and how they change across the latent space.
A.1.3 Optimizing locality placement
For locality placement, we try to match the relative in-degree centralities of samples when we look at the neighborhoods of all samples and the neighborhoods of localities (see Section 2 and Equation 2). We do this by optimizing parameters $\alpha$ and $\beta$ from Equation 2 using the $k$-means (here referred to as $\ell$-means) and DIRECT algorithms [36]. We use Manhattan instead of Euclidean distances to focus less on outliers and approximate centrality more equally across the samples (conceptually, this is similar to the difference between using mean absolute and mean squared errors). The rationale for using the weighing scheme in Equation 1 is forcing the computationally cheap $\ell$-means algorithm to optimize locality placement so that Equation 2 is minimized. We want to weigh samples based on their centrality, since the $\ell$-means algorithm optimizes Euclidean distances rather than our objective. Furthermore, the weights in Equation 1 factor in the inverse average distances to offset the fact that $\ell$-means optimizes the sum of squared Euclidean distances between points, which means that lower density regions of the latent space would have more probes (and localities), even when we weigh samples according to their centrality. Put together: we let the DIRECT algorithm optimize the weights of samples for the $\ell$-means algorithm, such that the localities produced by the $\ell$-means algorithm minimize our centrality-based loss function. Note that we use only the resulting centroids of the optimization as probes, while cluster assignments produced by the $\ell$-means algorithm are ignored. The $\ell$ localities are then simply the $n$-sized neighborhoods around each of the $\ell$ probes.

### v2

@@ 3 Experiments
We perform several experiments to verify and showcase the utility of LAVA explanations. First, we evaluate the stepwise and end-to-end stability of the LAVA method. Second, we analyze whether LAVA explanations yield the properties expected by design. Third, we compare LAVA against the only existing alternative, the label-free explainability framework (Crabbé and Schaar, 2022). Finally, we use LAVA to explain UMAP-embedded kidney single-cell gene expression data, and reveal patterns related to kidney disease.
@@
Datasets. Our experiments use two datasets: MNIST handwritten digit images (Lecun et al., 1998), and KPMP [^1] Kidney Precision Medicine Project: https://www.kpmp.org/. single-cell kidney gene expression (Lake et al., 2023). The KPMP dataset measures the activity of $\sim$30k genes (features) for $\sim$225k kidney cells (samples) (Stark et al., 2019). We remove mitochondrial and ribosomal genes, and use only the top $1000$ most variable genes, following single-cell analysis guidelines (Heumos et al., 2023) (Appendix B.2).
Experimental setup. We obtain two types of embeddings for MNIST, using UMAP manifold learning with default hyperparameters (McInnes et al., 2020)) or a denoising convolutional autoencoder (Crabbé and Schaar, 2022). For KPMP, we use the publicly available UMAP embeddings. We apply LAVA to each set of embeddings, with neighborhood size $n$ of $3000$ for MNIST and $500$ for KPMP (Appendix A.1), and upweighing factor $\nu=9$ for all experiments (Appendix A.3). To visualize MNIST results, we depict correlations of localities and modules as image-sized heatmaps, with pixel color denoting sum of correlations, and with blue lines between correlated pixels (Figure 1, Appendix B.4).
@@
Figure 4: LAVA explanations. Labeled UMAP embeddings and LAVA explanations (modules & presences) for a, MNIST and b, KPMP.
3.1 Stability of LAVA Explanations
We assess stepwise and end-to-end solutions of LAVA to verify its ability to deliver stable explanations. Setup. Applying LAVA to UMAP embeddings of MNIST and KPMP, we assess stability of locality placement, module extraction, and end-to-end explanations as follows. Locality placement: we run locality placement and representation (steps 1-2) $R_{L}=5$ times, for different values of overlap $o$; we then calculate $D\choose 2$ feature-pair average correlations across localities, and compare those between runs per $o$ value using cosine similarity (Figure 3a). AMF module extraction: using an arbitrarily chosen locality placement with $o=12$ for MNIST and $o=3$ for KPMP, whose $o$ values yield a compromise between fewer localities and higher stability, we run AMF module extraction (step 3) $R_{M}=10$ times for different numbers of modules $M$; we report reconstruction quality (main loss term as cosine similarity instead of distance), overestimation ratio, and silhouette width of AMF runs per $M$ (Figure 3b). End-to-end LAVA explanations, or combined locality placement & AMF module extraction: we run AMF for $M=9$ modules (criteria for $M$: Section 2.3) $R_{M}=10$ times for each of the $R_{L}=5$ locality placements; to assess similarity of the patterns underlying the modules irrespective of module organization (beyond silhouette width) across locality placements, we take a presence-weighted average of the module correlations per AMF run, average across runs per placement, and compare using cosine similarity (Figure 3c); for an arbitrarily selected placement (Figure 3c, highlighted), we also compare between presence-weighted module averages of individual AMF runs (Figure 3d).
LAVA locality placement is stable. Feature-pair correlations are highly similar across locality placements for both MNIST and KPMP (cosine similarity >0.993>0.993), even with smaller numbers of localities influenced by overlap $o$, indicating that locality placement is stable (Figure 3a, Appendix B.5). LAVA module extraction is stable. As expected, larger numbers of modules $M$ lead to better reconstruction of the underlying correlations (Figure 3b, red). For MNIST, we find relatively better reconstruction with small standard deviations (Figure 3b, red), but also larger instability denoted by smaller silhouette widths (Figure 3b, blue), pointing to successful optimization with multiple valid solutions. Here we note that silhouette width focuses on the goal of identifying a similar modular organization, which goes beyond capturing similar underlying patterns and could be too strict an objective for complex data. We therefore also evaluate pattern similarity below, and in Appendix B.7. For KPMP, a decrease in silhouette widths with increasing $M$ (Figure  3b, blue) indicates that smaller numbers of modules $M$ result in higher stability. Just like for MNIST, similar trends in reconstruction and stability for KPMP suggest the existence of multiple valid solutions. LAVA delivers end-to-end stable explanations. The LAVA modules capture highly similar correlation patterns across the $R_{L}=5$ locality placements (Figure 3c), and across AMF runs of a locality placement (Figure 3d). This indicates that LAVA explanations are highly stable end-to-end, throughout steps 1 to 3 of the LAVA method. Together, high reconstruction quality with small standard deviations and highly similar patterns (Figures 3b-3d), show that module runs capture very similar underlying patterns. We therefore conclude that the module runs represent differently factorized but equally valid solutions.
@@ 3.1 Stability of LAVA Explanations
Figure 5: LAVA module co-presences and locality reconstruction. Co-presence as entropy of module presences per locality for UMAP embeddings of a, MNIST and b, KPMP. c, Reconstruction of an example MNIST locality using its co-present modules.
3.2 Properties of LAVA Explanations
We investigate if the explanations extracted by LAVA possess the properties sought after by design. Setup. We examine the LAVA modules obtained for UMAP embeddings of MNIST and KPMP as described in Section 3.1.
LAVA modules reveal globally reoccurring local subpatterns. For both datasets, the extracted modules represent globally reoccurring subpatterns not tied to neighboring localities (Figure 4). For instance, for MNIST, the two modules with the highest sum of presence in the latent space capture correlation patterns shared between digits $3$, 55, and $8$, and between digits $4$, 77, and $9$ (Figure 4a). We observe similar trends for KPMP embeddings (Figure 4b), with some modules present across most major cell types (module $3$), and others pertaining to more local regions of specific cell types (e.g. epithelial cells for modules $1$, 22, and $4$). The module extraction uses no information about the spatial organization of localities, allowing it to identify subpatterns shared across or varying within clusters. LAVA modules identify reoccurring, not averaged subpatterns. Modules exhibit overestimation error ratios below 15% (Figure 3b). Errors made up of predominantly underestimation indicate that locality-averaged patterns are avoided in favor of sparser but highly present patterns, much like in quantile regression (Koenker and Hallock, 2001). Combined with the semantics of cosine similarity, the presence of a module in two localities implies not only a shared correlation subpattern, but also that such a subpattern can directly be observed in each of the localities (Figure 4, Appendices B.7-B.8). We further compare LAVA modules, extracted using AMF, to NMF components (Appendix B.8). Without control for overestimation, NMF yields ratios of 50%, indicating that NMF components capture average patterns across smaller sets of localities rather than finding more globally reoccurring subpatterns like AMF. Moreover, NMF relies on weighted sums of components, which is incoherent for correlations. Multiple LAVA modules can coexist in a locality. To investigate co-presence of LAVA modules, we scale their presences to sum to $1$ per locality and calculate Shannon entropy. Ignoring localities with non-scaled presence sums <0.5<0.5 to prevent inflated entropies from scaling, we get an average locality entropy of $0.68$ ($\pm 0.41$) for MNIST and $1.41$ ($\pm 0.75$) for KPMP (Figures 5a-5b). For reference, 0 means no module co-presence and $3.17$ corresponds to equal presence of all $9$ modules. The obtained entropies indicate module co-presences in both datasets, and more so in KPMP (Figures 5a-5b; Figure 5c co-presence in a locality). Allowing co-presence, a locality can be related to different subsets of localities via presences of modules, which may represent partially overlapping subpatterns of correlation.
Figure 6: LAVA & label-free XAI. Example (left) LFXAI feature importances and (right) LAVA explanations – modules and presences – for localities of MNIST autoencoder embeddings.
3.3 Comparing LAVA and Label-Free Explainability
We perform a qualitative comparison between LAVA and the label-free explainability framework (LFXAI) by Crabbé and Schaar (2022), to our knowledge the only other explainability method for unsupervised latent spaces. Comparison is challenging, since LAVA and LFXAI are fundamentally different. LAVA is natively designed to explain unsupervised models, while LFXAI uses adaptations of XAI for supervised ML. We focus on LFXAI-based feature importance, as it relates original features to embeddings, and is thus conceptually closer to LAVA than example importance. In this context, LFXAI explanations are per sample linear combinations of feature importances derived independently for each latent dimension as model output. By contrast, LAVA captures feature covariation subpatterns across neighborhoods of sample embeddings in the multidimensional latent space. Setup. We generate $4$-dimensional MNIST embeddings with a denoising convolutional autoencoder, using the implementation of Crabbé and Schaar (2022). Note that LFXAI does not support embeddings without mapping function like UMAP. For LFXAI explanations, we use LFXAI with gradSHAP (Lundberg and Lee, 2017) to generate feature importances for every sample. For LAVA explanations, we run LAVA with number of neighbors $n=3000$, overlap $o=12$, and number of modules $M=8$ (Appendix B.7.3).
Explaining samples vs. embedding structure. To enable comparison, we analyze explanations at the level of LAVA localities. We average LFXAI feature importances of the samples in each locality, and compare to LAVA modules for the same locality (Figure 6, Appendix B.7.4).
@@ 3.3 Comparing LAVA and Label-Free Explainability
Figure 7: LAVA module 4 for KPMP. Module presence, chronic kidney disease (CKD) percentage, module-CKD overlap, gene analysis.
The LFXAI locality-averaged importances resemble the original digit images (Figure 6, LFXAI), consistent with the purpose of revealing which features were relevant to place a sample in the latent space relative to a reference, empty image. However, it remains nontrivial if and how LFXAI explanations can be exploited to understand the structure of the embeddings, regardless of which feature importance method is used. In contrast, LAVA explains how input features relate to the structure of the embeddings at a desired level of granularity. By capturing reoccurring feature subpatterns across localities, LAVA can further reveal relations between seemingly distant regions of the latent space that would not be apparent otherwise (Figure 6, LAVA). LAVA enables discovery beyond explanation. The LAVA and LFXAI methods provide complementary explanations for latent embeddings, each from a different perspective. Since one key goal of unsupervised learning is to identify patterns leading to new discoveries, the ability of LAVA to uncover hidden structure makes it ideally positioned to drive knowledge discovery beyond explainability (Figures 4 and 6). Importantly, when relating features to embedding structure, LAVA detects feature covariation irrespective of feature usage by the model. This allows LAVA to include features otherwise ignored, for instance for exhibiting strong dependencies with others. As a result, LAVA delivers more complete explanations that address a common limitation of explainability methods (Molnar, 2022), and circumvent the Rashomon effect (Breiman, 2001; Rudin et al., 2024).
3.4 Explaining gene expression embeddings
Finally, we enquire if LAVA can reveal biologically relevant patterns in UMAP embeddings of KPMP data. Setup. We analyze LAVA modules for KPMP (Section 3.1), together with metadata on the human donors of the cells: we look at disease state, which is either healthy, acute kidney failure (ACF), or chronic kidney disease (CKD). To quantify the association between modules and a given disease state, we calculate Pearson correlation and respective p-value between module presence and percentage of samples labeled with the disease state (ignoring localities where a module is not present). We further investigate if genes in those modules are involved in biological pathways related to the diseased state (Appendix B.9).
Module correlates with diseased samples. Cells from donors with different disease states are spread across the latent space, with no apparent clustering. However, presence of module $4$ is positively correlated with the percentage of diseased samples ($0.23$, $p<0.01$), and specifically CKD (Figure 7, $0.28$, $p<0.01$). Together with the fact that module $4$ is present in epithelial cells and not blood cells (Figures 4b and 7), this suggests that module $4$ could be biologically relevant for kidney disease and CKD-related pathways. Module genes and disease related pathways. Overrepresentation analysis of biological pathways (Fang et al., 2022) for the genes in module $4$ identifies $21$ significant pathways ($p<0.01$). To highlight biological relevance, we single out the “cellular response to stress” and “DNA damage response” pathways, linked to CKD in previous studies (Patera et al., 2024; Schupp et al., 2016; Molitoris, 2019; Wang et al., 2023b). Overall, our findings suggest that the local organization of KPMP latent UMAP embeddings across kidney cell types exhibits patterns of gene correlation related to kidney disease. This local structure is not apparent in the clustering of the data, but is notably discoverable through the LAVA modules.
@@ 4 Conclusions
In this work, we frame the problem of explainability of latent embeddings as enabling reasoning about the latent spatial organization through the underlying input data. To that effect, we propose the LAVA method, which generates explanations in the form of modules, capturing local subpatterns of feature correlation reoccurring globally throughout the latent space. We show that LAVA produces stable explanations at a desired level of granularity, complementary to existing label-free explainability approaches. In addition, LAVA uncovers hidden structure not apparent otherwise, which makes it especially suited for driving knowledge discovery. Specifically, we show that LAVA reveals domain-relevant patterns, such as visual parts of images and disease signals in cellular processes. For future work, we envision further exploration of spatially-guided reasoning about latent embeddings, especially for higher-dimensional latent spaces and supervised learning.
Acknowledgments
@@
The Kidney Precision Medicine Project (KPMP) is supported by the National Institute of Diabetes and Digestive and Kidney Diseases (NIDDK) through the following grants: U01DK133081, U01DK133091, U01DK133092, U01DK133093, U01DK133095, U01DK133097, U01DK114866, U01DK114908, U01DK133090, U01DK133113, U01DK133766, U01DK133768, U01DK114907, U01DK114920, U01DK114923, U01DK114933, U24DK114886, UH3DK114926, UH3DK114861, UH3DK114915, and UH3DK114937. We gratefully acknowledge the essential contributions of our patient participants and the support of the American public through their tax dollars.
The authors acknowledge the University of Michigan Medical School Central Biorepository (RRID:SCR_026845) for providing biospecimen storage, management, and distribution services in support of the research reported in this publication/grant application/presentation.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ A.1 Locality Definition
A.1.1 Defining the Size of Localities
@@
The choice of hyperparameter $n$ is important as changing the size of localities can have a substantial influence on correlations that LAVA will calculate in step (2), and subsequently for module extraction in step (3). A data-driven way of choosing the value $n$ can be looking at which size of latent neighborhood most resembles the neighborhoods in the original feature size. In choosing such an $n$, we attempt to capture localities at a scale that most accurately represents organization in the original feature space. Another, more targeted, approach is motivating the choice of $n$ using metadata of interest. This might ground the scale of localities to existing knowledge and make findings more relevant for a specific downstream task. We employ both approaches in our experiments in Section 3, and deliberate further in B.3.
A.1.2 Optimization of Locality Placement
For locality placement, we minimize the difference between the relative in-degree centralities of samples when we look at the neighborhoods of all samples, and the neighborhoods of localities (see Section 2 and Equation 2). For example, if some sample is in the $n$-nearest neighborhoods of $\%1$ of the samples in the embedding space, we want our localities to be placed so that the sample is a part of $\%1$ of the localities, i.e., so that the sample retains its in-degree centrality. We do this by optimizing parameters $\alpha$ and $\beta$ from Equation 2, utilizing the $k$-means (here referred to as $\ell$-means) and DIRECT algorithms (Jones et al., 1993).